### v1

# Application of Whisper in Clinical Practice: the Post-Stroke Speech Assessment during a Naming Task thanks:

Abstract
Detailed assessment of language impairment following stroke remains a cognitively complex and clinician-intensive task, limiting timely and scalable diagnosis. Automatic Speech Recognition (ASR) foundation models offer a promising pathway to augment human evaluation through intelligent systems, but their effectiveness in the context of speech and language impairment remains uncertain. In this study, we evaluate whether Whisper, a state-of-the-art ASR foundation model, can be applied to transcribe and analyze speech from patients with stroke during a commonly used picture-naming task. We assess both verbatim transcription accuracy and the model’s ability to support downstream prediction of language function, which has major implications for outcomes after stroke. Our results show that the baseline Whisper model performs poorly on single-word speech utterances. Nevertheless, fine-tuning Whisper significantly improves transcription accuracy (reducing Word Error Rate by 87.72% in healthy speech and 71.22% in speech from patients). Further, learned representations from the model enable accurate prediction of speech quality (average F1 Macro of 0.74 for healthy, 0.75 for patients). However, evaluations on an unseen (TORGO) dataset reveal limited generalizability, highlighting the inability of Whisper to perform zero-shot transcription of single-word utterances on out-of-domain clinical speech and emphasizing the need to adapt models to specific clinical populations. While challenges remain in cross-domain generalization, these findings highlight the potential of foundation models, when appropriately fine-tuned, to advance automated speech and language assessment and rehabilitation for stroke-related impairments.
Index Terms:
Automatic speech recognition, Clinical speech analysis, Language impairment detection
I Introduction
Stroke is one of the leading causes of adult death and disability worldwide, with global incidence of 12.2 million per year [1]. One of the most debilitating and common consequences after stroke is aphasia, an acquired disorder of speech and language production and comprehension, with over 30% prevalence in patients with stroke[2, 3, 4]. Due to high heterogeneity of aphasic symptoms, diagnosis requires an in-depth assessment to uncover specific impairments and tailor speech therapy to individual patient’s needs. Computerized assessments have the advantage, compared to standard pen and paper tests, of being administered with minimal supervision, providing a cost-effective and accessible platform for speech and language evaluation of patients with stroke in varied settings [5]. However, these digital assessments often depend on manual speech transcription and evaluation by a trained clinician, which is time-consuming, prone to inconsistencies due to inter- and intra-rater variability and are difficult to implement on a large scale.
Advancements in automatic speech recognition (ASR) methods can fill this gap and allow for automation of speech evaluation pipelines. However, whilst the latest ASR models perform exceedingly well on healthy speech data [8], several limitations have hindered the translation of ASR to pathological speech analysis. Firstly, currently available ASR models have been trained on healthy speech, limiting their generalizability to pathological speech. Secondly, the potential variability in speech impairments in aphasia necessitates large clinical speech databases for effective training of ASR models, which are currently relatively limited.
The extensive linguistic knowledge embedded in foundation ASR model, acquired through pretraining on large and diverse datasets, has the potential to be leveraged in a zero and few shot learning context to automate post-stroke speech evaluation pipelines [6, 7].
In this study, we aimed to assess the performance of Whisper [8], a state-of-the-art transformer-based ASR model pretrained on 680,000 hours of multilingual data, on a stroke-specific speech database derived from a Naming task, commonly used in post-stroke speech assessments. We fine-tuned Whisper on speech from age-matched healthy older adults and patients with stroke and compared the verbatim transcription performance of the fine-tuned and the baseline Whisper models. Further, we evaluated whether Whisper-derived representations could be effectively applied to clinically-relevant downstream tasks, specifically speech impairment severity prediction. Lastly, we assessed the quality of these predictions by conducting a divergent and convergent validity analysis with clinical stroke outcomes.
II Related Work
Current approaches for automated speech impairment analysis involve three key steps. First, speech-to-text ASR models are employed to generate speech transcriptions from audio recordings. Then, a feature extraction process is applied to the produced transcriptions and audio data. The extracted lexico-acoustic features are used in downstream analyses, such as classification and regression, to estimate the degree of speech impairment [9]. Different algorithms have been used to complete each one of these steps.
II-A Recurrent Neural Networks
Previous studies have trained customized ASR models to transcribe aphasic speech, subsequently deriving acoustic and transcription-based features to estimate aphasia severity [11, 10]. Le and colleagues[11] utilized a Bidirectional Long-Short Memory-Recurrent Neural Network (BLSTM-RNN) model trained on acoustic features of aphasic speech derived from the AphasiaBank dataset [12] and frame-level senone and monophone labels, obtained through a Hidden Markov Model-Gaussian Mixture Model (HMM-GMM) system, to produce speech transcriptions. Similarly, Qin and colleagues[10] utilized a Time-Delay Neural Network combined with a Bidirectional Long-Short Term Memory (TDNN-BLSTM) model to transcribe Cantonese aphasic speech. The model was trained on acoustic features derived from healthy speech and HMM-GMM-derived triphone target labels [10]. The ASR transcriptions were used to extract features based on text alone [11], or in combination with audio features [10] to estimate aphasia severity by regression. While the above studies reached promising accuracy, the employed ASR systems had complex architectures, which required a multi-step training process, limiting global optimization [13]. Additionally, these models were based on variants of RNN architectures, which are constrained in modeling long-range dependencies compared to Transformers [14].
II-B Transformers
The advent of Transformers has revolutionized the field of ASR by implementing self-attention mechanisms, which capture temporal relationships across the entire input sequence simultaneously, thus overcoming the range limitations of RNN-based architectures and enabling more efficient encoding of contextual information [15, 16]. Multiple ASR transformer-based models are currently available for speech-to-text transcription including Transformer Transducer [17], wav2vec 2.0 [18], SpeechStew [19], Conformer [20] and Whisper [8]. However, while transformer-based models have demonstrated exceptional performance on healthy speech ASR [8], they often encounter challenges in adapting to aphasic speech recognition due to the high heterogeneity in the lexico-acoustic features of post-stroke speech.
Recent studies have leveraged pretrained transformer-based ASR models, adapting them to enhance performance on aphasic speech. Within this framework[21], XLSR-53, a model built upon wav2vec 2.0 architecture, was fine-tuned to adapt it to aphasic speech transcription. XLSR-53 [22], pretrained on the extensive dataset, which contains 56,000 hours of unlabeled audio, uses convolutional neural network layers to generate latent speech representations [18]. These representations are then passed through a transformer network to encode contextualized speech features [18]. However, wav2vec 2.0 architecture is outperformed by other transformer-based foundation models such as Whisper [8, 23]. Whisper-derived representations have been successfully utilized in previous studies to classify dysfluencies in stuttering [24] and to estimate dysarthria severity [25]. Further, Whisper and XLSR-53 model transcripts have been successfully used for relevant speech feature extraction, later used for aphasia type classification, reaching average 90.6 F1 score [26]. Recent work also investigated hyperparameter tuning and fine-tuning of Whisper to better fit aphasic speech, reaching promising results of 38.5% and 21.93% WER on aphasic speech transcription, respectively. [27, 28]. However, the vast majority of previous work has focused on sentential speech, and the applications of Whisper to single-word utterances remain to be studied.
III Method
Figure 1: General study pipeline. Overview of the training and inference pipeline for verbatim transcription and accuracy prediction models. Transcription: Training – Whisper models were fine-tuned on speech data from 5 different datasets. A synthetic dataset of audio-transcription word pairs based on stimuli in the Naming task, SONIVA-Naming healthy dataset, collected from healthy age-matched participants performing the Naming task, SONIVA-Naming patient dataset, collected from patients with stroke performing the Naming task, a combined dataset of SONIVA-Naming healthy and patient data, and a combined dataset of all available data (synthetic, SONIVA-Naming healthy and SONIVA-Naming patient). All models were trained using cross-entropy loss to transcribe spoken words (e.g. “acorn”) from the Naming task input audio. Inference – Fine-tuned models (ft-syn, ft-h, ft-p, ft-hp and ft-all) were evaluated on 4 separate datasets – synthetic, SONIVA-Naming healthy, SONIVA-Naming patient, and an unseen testing dataset of dysarthric speech derived from the TORGO database. Accuracy: Training – accuracy prediction models were trained on SONIVA-Naming derived healthy and patient speech in a linear probing framework. The encoder layer of the models was frozen and the classification projection head was trained to predict accuracy scores using cross-entropy loss (e.g “2”). The encoder weights of the models were initialized at baseline or with the weights derived from the models trained for verbatim transcription of healthy or patient speech (ft-h or ft-p).
III-A Dataset
Speech data from age-matched controls and patients with stroke were derived from participants undergoing speech and cognitive testing as part of the the IC3 study [5]. Specifically the data was obtained from the Naming task, modified from the 30-item Boston Naming Task [29], where participants were instructed to name a depicted black and white line drawing picture with a single word. The speech recordings from this task, amongst other speech production tasks, have contributed to a larger speech database SONIVA (Speech Recognition Validation in Aphasia) [30]. The data used in this study will be referred to as SONIVA-Naming henceforth. The accuracy of the Naming task performance was evaluated across each trial using four accuracy metrics: semantic content (Semantic), fluency of articulation (Dysfluency), presence of self-corrections in the response (Self-correction) and phonological correctness (Phonology). Each trial was scored by trained raters on a 3-point scale: 0, 1, or 2, where 2 corresponds to the highest accuracy. The verbatim transcriptions of the audio files ranged from single word to sentences depending on the participants responses. The inter-rater reliability against a qualified speech therapist was high (Mean Intraclass Correlation Coefficient = 0.83$\pm$ 0.0046), and all ambiguous cases were resolved by consulting a speech therapist. In total, the SONIVA-Naming healthy dataset comprised 3960 Naming task trials from 132 healthy participants, while the SONIVA-Naming patient dataset consisted of 2609 trials from 87 patients (see Table I). There was no group difference for age (Mann-Whitney U statistic: $0.885542.50,p=$ 0.88). Patients had lower education ($\chi^{2}$=26.19, $p<0.001$) and a higher proportion of non-native English speakers ($\chi^{2}$=4.02, $p=0.04$). A noise-free, single word, baseline synthetic dataset was additionally generated using the Google Text-to-Speech library for Python [31]. Seven English accents were used to generate a total of 210 synthetic audio files across the 30 stimuli words available in the Naming task.
An additional unseen testing dataset was used for assessment of model generalizability for single-word pathological speech transcription. This dataset was derived from the TORGO database [32]. The database includes speech recordings from individuals with dysarthria, a disorder caused by disruptions in the speech motor system, common in conditions such as cerebral palsy and amyotropic lateral sclerosis. The single-word recordings from the database were used to assess model transcription of short-form data. The TORGO dataset consisted of data from 7 participants (4 male and 3 female) and comprised a total of 1240 utterances.
TABLE I: Sample demographics of the SONIVA-Naming dataset
Demographics feature SONIVA-Naming healthy SONIVA-Naming patients
(N = 132) (N = 87)
Mean (Standard Deviation)
Age 61.6 (10.8) 61.8 (13.6)
Sex***
     Male:Female 52:77 62:25
     Missing data 3 0
English Language*
      Non-native 27 30
      Native 102 57
Education level***
     School 36 45
     Degree 43 37
     Post-graduate 37 3
     Missing data 16 2
${}^{\mathrm{**}}\;p<0.05$; ${}^{***}\,p<0.001$
III-B Data Pre-processing
All audio data were resampled to 16 kHz. In the main dataset used for model training and evaluation (SONIVA-Naming), all audio recordings without at least one ground truth accuracy metric (Semantic, Dysfluency, Self-correction and Phonology) were excluded when fine-tuning models for the accuracy score prediction task ($n=23$ and $n=114$ for SONIVA-Naming healthy and SONIVA-Naming patient datasets, respectively). The data were divided into training, validation and test sets in the ratio of 7:1:2. The splitting was completed at the participant level such that all trials from the same participants were in the same partition. Due to the high class imbalance across the accuracy scores, the training sets for the accuracy prediction task were randomly downsampled to the minority class in each accuracy metric. This resulted in 144, 204, 90 and 21 audio files for the SONIVA-Naming healthy dataset for Semantic, Dysfluency, Self-correction and Phonology, respectively. In case of the SONIVA-Naming patient dataset, the final sample size was 198, 729, 108, and 171 trials for the same metrics respectively.
III-C Whisper Architecture
Whisper has an encoder-decoder Transformer architecture [8] with different model size availability, corresponding to the number of parameters (39M-1550M) and encoder-decoder block layers (4-32). Each encoder block of Whisper consists of a self-attention layer and a multilayer perceptron. The model processes audio input in a form of an 80-channel log-Mel spectrogram, passing it through two initial convolutional layers followed by the encoder blocks. The encoder outputs information-dense context vectors, which are a high-level feature representation of the input audio sequence, encompassing acoustic, positional, and contextual information through self-attention mechanisms and positional embedding.
The decoder blocks mirror the architecture of the encoder blocks, incorporating an additional cross-attention layer, which enables the decoder to focus on the outputs of the encoder. This architecture allows the decoder to process the context vectors produced by the encoder and the tokenized input text sequences. Whisper utilizes a Byte-Pair Encoding tokenizer to break down the input text into smaller sub-word units [33]. During training, tokenized text sequences correspond to the audio input’s ground truth labels, while during inference, they consist of previously predicted tokens. Additionally, Whisper utilizes a range of special tokens, which signal the start and end of the sequence, task type and language. The decoder’s raw output is passed through a final softmax layer to obtain the next token prediction.
To complete the accuracy classification task, the same encoder architecture was used, but the decoder was replaced with a linear classifier projection head, consisting of two linear layers.
III-D Model Training
Verbatim Transcription Task
Whisper was separately fine-tuned on the training set of synthetic, SONIVA-Naming healthy, and SONIVA-Naming patient data, as well as combined datasets of all participant data (SONIVA-Naming healthy and SONIVA-Naming patient) and all available data (synthetic, SONIVA-Naming healthy and SONIVA-Naming patient). Small and Medium Whisper models were used to determine the optimal model size. In total, ten fine-tuned models were obtained: fine-tuned on synthetic data (ft-syn), on SONIVA-Naming healthy data (ft-h), on SONIVA-Naming patient data (ft-p), on all participant data (ft-hp) and all available data (ft-all), each in Small and Medium sizes.
Models were trained for a maximum of 1000 steps, with a batch size of 16. A learning rate of $1\times 10^{-5}$ was used, with a linear learning rate scheduler incorporating 250 warm up steps. The model parameters were updated with AdamW optimiser using Cross-Entropy loss function (1).
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\log(\hat{y}_{ij}),$ (1)
where N is the total number of samples, C is the total number of classes, $y_{ij}$ is the true label and $\hat{y}_{ij}$ is the predicted probability for class $j$ for instance $i$.
Training was conducted on one NVIDIA RTX 6000 GPU. Models were evaluated on the validation set (batch size of 8) at 50 step intervals. The best model was chosen based on Word Error Rate (WER) on the validation set and was used for final evaluation on the test sets. The evaluation was completed on four different test sets, namely the synthetic, SONIVA-Naming healthy, and SONIVA-Naming patient as well as the unseen TORGO database.
Accuracy Prediction Task
The setup was formulated as a multi-class classification task, where one of three possible accuracy scores (0, 1 or 2) were predicted for each given Naming task trial recording in the SONIVA-Naming healthy and patient datasets.
The encoder was initialized either with the baseline pretrained encoder weights (Whisper without fine-tuning) or with the encoder weights of the models fine-tuned for verbatim transcription of healthy and patient speech. The ft-h and ft-p model weights were used when predicting accuracy scores on SONIVA-Naming healthy and SONIVA-Naming patient data, respectively, to capture the most relevant speech representations of each group. This setup resulted in a total of 32 models: models were trained separately for each accuracy metric (Semantic, Dysfluency, Self-correction, Phonology), with two encoder weight configurations (fine-tuned or baseline), each in two sizes (Small and Medium) for each dataset separately (SONIVA-Naming healthy and SONIVA-Naming patient). During training, the encoder weights were frozen and only the linear classification head was trained (i.e., linear probing).
Models were trained for a maximum of 8000 steps and with a batch size of 16. A learning rate of $1\times 10^{-5}$ was used, with a linear learning rate scheduler incorporating 500 warm up steps. Models were evaluated on the validation set at 100 step intervals. The model parameters were updated with AdamW optimizer using Cross-Entropy loss function (1). The best model was chosen based on F1 Macro metric on the validation set and was used for final evaluation on the test set.
III-E Evaluation metrics
Word Error Rate (WER)
The WER was defined based on the string edit distance, which consists of the ratio of necessary string modifications needed to convert the model prediction into the ground truth label divided by the number of total words spoken (2).
${WER}=\frac{S+I+D}{N}$ (2)
where S is the number of substitutions, I of insertions, D of deletions, and N is the total number of words spoken. The score was then multiplied by 100 to obtain a percentage measure.
F1 Macro
Performance on the accuracy score prediction task was assessed based on binarized F1 scores. Specifically, the accuracy score 1 was merged with 0 to derive an impaired class (class 0), while accuracy score of 2 was converted to 1 (class 1) to represent unimpaired responses. The F1 score for the impaired and unimpaired classes were calculated, and then the average between the two was used to obtain the F1 macro, which was used as the main performance metric (3).
${F1_{Macro}}=\frac{\sum_{i=1}^{n}F1_{i}}{n}$ (3)
where n is the number of classes.
Target Word Detection Accuracy
The verbatim transcriptions of the SONIVA-Naming dataset trials often consisted of more than the target Naming task word. A target word detection accuracy was calculated by assessing whether the predicted transcriptions correctly identified the target word exclusively. A prediction was classified as True Positive when the target word was present in both the ground truth transcription and prediction, and True Negative when it was absent in both. The presence of the target word in the ground truth label but not in the prediction indicated a False Negative and the reverse - a False Positive.
III-F Statistical Analysis
Model size and type analysis
The effect of model type (i.e., baseline vs fine-tuned) and size (i.e., small vs medium) on WER across all trials was assessed on the test set using the Friedman test, followed by Mann-Whitney U tests with Bonferroni correction for multiple comparisons. The non-parametric nature of the tests was due to the non-normal distributions of the WER within each dataset.
Predictive Validity Analysis
To evaluate the clinical validity of the predicted speech accuracy scores, an overall predicted accuracy score was derived for each patient in the SONIVA-Naming patient test set ($n=18$) by averaging their predicted trial-by-trial scores for each metric. Patients were classified as either impaired or unimpaired on each predicted accuracy metric. Patients with an overall predicted score smaller or equal to 1 were classified as impaired. Otherwise, they were classified as unimpaired. Impairment status was validated against multiple known clinical and demographic factors. Convergent validity was tested against a manually assessed speech fluency metric, stroke history and English as a second language status – hypothesized to affect predicted impairment status. Divergent validity was tested against sex, low density lipoprotein (LDL) cholesterol levels and smoking status, hypothesized to not show any relationship with predicted impairment status.
To complete the analysis, all categorical variables (e.g., Sex, Smoking status, English as second language, Previous Stroke), were one-hot encoded. The distribution of continuous variables (LDL Cholesterol, speech fluency) was tested for normality using the Shapiro-Wilk test. Since the assumption of normality was met, continuous variables between groups were compared with a Student’s t-test. Categorical variables were compared between groups using Fisher’s exact test. Statistical significance was set at $p<0.05$.
IV Results
IV-A Verbatim Transcription
A significant effect of model size and type on WER was detected when testing on synthetic ($\chi^{2}$(11, $N$=42) = 313.95, $p<0.001$), SONIVA-Naming healthy ($\chi^{2}$(11, $N$=806) = 5215.60, $p<0.01$), and SONIVA-Naming patient datasets ($\chi^{2}$(11, $N$=524) = 2442.80, $p<0.01$). All fine-tuned models significantly outperformed baseline Whisper in transcription accuracy on the synthetic, healthy and patient datasets ($p<0.01$) (Table II). Fine-tuning on healthy and patient speech resulted in 87.72% and 71.22% improvement in WER for healthy and patient speech compared to baseline model of the same size. A significant improvement was also observed in case of synthetic data, where the best-performing models were Medium ft-h, Medium ft-p and Small and Medium ft-all, reaching a WER of 0%, compared to a WER of minimum 85.71% in case of Whisper baseline. The best-performing model on the healthy data was the Small ft-h, reaching average WER of 8.82%. On the patient dataset, Medium ft-hp reached the lowest average WER of 26.35%. However, this Medium model was not significantly better than its Small counterpart ($p>0.05$). Additionally, when evaluated on the patient dataset, the difference in performance between ft-hp and single-dataset trained models ft-h and ft-p was insignificant ($p>0.05$) in both model sizes.
On the unseen TORGO database, a significant effect of model size and type on WER was also detected ($\chi^{2}$(11, $N$=1240) = 850.08, $p<0.001$) (Table III). All fine-tuned models significantly outperformed baseline Whisper in verbatim transcription accuracy ($p<0.01$). The best performing model was Medium ft-syn, which achieved a 22.64% reduction in WER compared to the baseline model in the same size. However, it did not perform significantly better than other fine-tuned models ($p>0.5$), with the exception of Medium ft-hp ($p=0.045$).
TABLE II: Comparison of word error rate in transcription across testing datasets of SONIVA-Naming data.
Word Error Rate (%)
Size
Testing Dataset Model Small Medium
Synthetic data Baseline 92.85 85.71
ft-syn 45.23 21.42
ft-h 2.38 0
ft-p 7.14 0
ft-hp 2.38 2.38
ft-all 0 0
SONIVA-Naming healthy Baseline 96.54 90.87
ft-syn 53.64 48.06
ft-h 8.82 9.89
ft-p 14.60 11.94
ft-hp 12.69 12.58
ft-all 12.00 9.07
SONIVA-Naming patient Baseline 100.54 97.57
ft-syn 75.83 71.53
ft-h 28.38 27.72
ft-p 28.82 26.80
ft-hp 27.79 26.35
ft-all 27.04 29.15
TABLE III: Comparison of word error rate in transcription on the out-of-domain TORGO database
Word Error Rate (%)
Size
Testing Dataset Model Small Medium
TORGO database Baseline 97.82 96.13
ft-syn 77.72 73.49
ft-h 77.39 75.65
ft-p 75.01 74.62
ft-hp 78.92 79.67
ft-all 78.10 76.58
IV-B Target Word Detection accuracy
Compared to baseline Whisper, fine-tuned models had a better performance on target word detection accuracy in both healthy and patient speech (Table IV). For healthy speech, the best-performing model was the Small ft-h, reaching an accuracy of 0.97. This increase in performance was due to the ft-h model making noticeably fewer False Negative mistakes ($n=27$) compared to baseline Whisper ($n=496$). On the other hand, for patient speech, the best-performing models were the Small and Medium ft-hp, reaching an accuracy of 0.92. This model made fewer False Negative mistakes ($n=39$) than the Medium baseline model ($n=286$). However, Small and Medium ft-hp models produced 2 False Positive mistakes, while the Small and Medium baseline Whisper only had 1 and 0 False Positive mistakes, respectively. Overall, an improvement of up to 61% and 51% was observed for healthy and patient data, respectively.
TABLE IV: Model target word detection performance on the SONIVA-Naming dataset
Testing Dataset Size Model Accuracy True Positive True Negative False Positive False Negative
SONIVA-Naming healthy Small Baseline 0.36 256 24 0 496
ft-syn 0.60 442 24 0 310
ft-h 0.97 725 24 0 27
ft-p 0.92 693 24 0 59
ft-hp 0.96 723 24 0 29
ft-all 0.96 719 24 00 33
Medium Baseline 0.40 284 24 0 468
ft-syn 0.59 432 24 0 320
ft-h 0.96 724 24 0 28
ft-p 0.95 716 24 0 36
ft-hp 0.95 714 24 0 38
ft-all 0.96 721 24 00 38
SONIVA-Naming patient Small Baseline 0.41 126 87 1 310
ft-syn 0.52 188 87 1 248
ft-h 0.91 390 85 3 46
ft-p 0.90 383 86 2 53
ft-hp 0.92 397 86 2 39
ft-all 0.90 389 85 3 47
Medium Baseline 0.45 150 88 0 286
ft-syn 0.48 165 87 1 271
ft-h 0.91 393 85 3 43
ft-p 0.91 389 86 2 47
ft-hp 0.92 397 86 2 39
ft-all 0.88 375 86 2 61
IV-C Accuracy Score Prediction
The best-performing models for predicting the Semantic, Dysfluency, and Self-correction accuracy scores during the Naming task on healthy data were the Medium models initialized with ft-h weights, achieving F1 Macro scores of 0.7449, 0.8390, and 0.7539, respectively (Table V). For the Phonology metric, the best-performing model was the Small model also initialized with ft-h weights, reaching an F1 Macro of 0.6424. Compared to their respective baselines, an average improvement of 7.24 ± 4.01% was observed across all accuracy metrics. Although all models had moderate to high F1 Macro scores, there was a discrepancy between their performance on correct/unimpaired (class 1) and incorrect/impaired (class 0) trials. The F1 scores for class 1 trials were 0.9814, 0.9564, 0.9860, and 0.9744, whereas the F1 scores for class 0 trials were 0.5085, 0.7215, 0.5217, and 0.3103 for the Semantic, Dysfluency, Phonology, and Self-correction metric models, respectively.
In context of patient speech, the best-performing model for predicting Semantic accuracy scores was the Small model initialised with ft-p encoder weights, reaching F1 Macro of 0.7659. For predicting Dysfluency and Self-correction, the best-performing models were the Small models initialized with baseline encoder weights, reaching F1 Macro of 0.9021 and 0.7112, respectively (Table V). For Phonology, the best model was the Medium model initialized with ft-p encoder weights, achieving F1 Macro of 0.6435. The observed improvement of fine-tuned models in Phonology and Semantics was in average 4.425 ± 0.175%. In Dysfluency and Self-correction, the baseline models performed on average 1.45±0.06% better than corresponding models initialized with ft-p encoder weights. While the discrepancy between performance on class 1 and class 0 trials was still evident, it was lower than on healthy data. The F1 scores in class 1 trials were 0.9105, 0.9282, 0.8944 and 0.8604, whereas F1 scores in class 0 were 0.6213, 0.8759, 0.5279 and 0.4265 for Semantic, Dysfluency, Self-correction and Phonology metric models, respectively.
TABLE V: F1 Macro for Different Accuracy Metrics and Testing Datasets
Testing Dataset Size Encoder Weight Semantic Dysfluency Self-correction Phonology
SONIVA-Naming healthy Small Baseline 0.6196 0.7823 0.6124 0.5606
ft-h 0.6948 0.8375 0.6876 0.6424
Medium Baseline 0.7094 0.8009 0.6197 0.5334
ft-h 0.7449 0.8390 0.7539 0.6045
SONIVA-Naming patient Small Baseline 0.7234 0.9021 0.7112 0.5873
ft-p 0.7659 0.8870 0.6973 0.5771
Medium Baseline 0.7520 0.9017 0.6775 0.5975
ft-p 0.7370 0.8891 0.6613 0.6435
IV-D Predictive Validity Analysis
Self-correction accuracy scores were excluded from this analysis as only two patients were identified as impaired based on model predictions of this metric, making further statistical analysis impossible.
The predictions of the best-performing accuracy prediction models were used to identify patients as impaired or unimpaired on each metric (Small Whisper initialized with ft-p encoder weights, small Whisper initialized with baseline encoder weights, and Medium Whisper initialized with ft-p encoder weights for Semantic, Dysfluency and Phonology, respectively).
Patients identified as impaired based on model predictions had significantly lower speech fluency than those identified as unimpaired based on Semantic ($t(16)=-3.61,p=0.01$), Dysfluency ($t(16)=-2.3880,p=0.02$) and Phonology ($t(16)=-2.3880,p=0.02$) metrics. Significantly more impaired patients spoke English as a second language compared to unimpaired patients, based on Dysfluency and Phonology metrics ($p=0.02$). Additionally, a higher number of patients with previous stroke history were identified as impaired than unimpaired based on Semantic ($p=0.02$), Dysfluency ($p=0.04$) and Phonology ($p=0.04$) metrics. The divergent validity analysis showed no significant differences between impaired and unimpaired patients for age, sex, LDL cholesterol or smoking status, as expected $(p>0.05)$.
Figure 2: Characteristics of patients identified as impaired (0) or unimpaired (1) from the SONIVA-Naming patient dataset. Predicted impairment status is displayed separately for Semantic (blue), Dysfluency (green), and Phonology (purple) metrics. Speech fluency, previous stroke history, English as second language (other languages), LDL cholesterol level, sex and smoking status were assessed. Error bars indicate standard deviation. Between group significance is denoted with p-values (uncorrected).
V Discussion
The performance of the baseline Whisper model during verbatim transcription of single-word Naming task speech in the healthy SONIVA-Naming dataset was significantly worse than previously reported performance on healthy continuous speech (Small Baseline Whisper WER 96.54% vs. 3.40%; Medium Baseline Whisper 90.87% vs. 2.90%) [8]. This suggests that baseline Whisper does not adapt directly to single-word speech data. Reassuringly, all fine-tuned Whisper models improved transcription performance across datasets, highlighting the necessity of fine-tuning. Models trained on healthy and patient data further improved the WER compared to those trained on synthetic data, suggesting the importance of natural speech diversity in improving transcription performance.
The fine-tuned model showed adequate performance for post-stroke speech transcription. In patients, the best performing model was the Medium ft-hp model achieving a 26.35% WER. This was not significantly different from the second-best performing model fine-tuned on patient data alone, Medium ft-p, (26.80% WER ; $p>0.05$), suggesting that fine-tuning on a combined patient and healthy dataset only negligibly improves performance.
The WER achieved by the ft-p model aligned with prior research on fine-tuned wav2vec 2.0 and Whisper models (WER 22.30%–55.50% depending on aphasia severity and 21.93% WER, respectively) [21, 30], and outperformed hyperparameter fine-tuned Whisper for continuous pathological speech (WER 38.50%) [27].
Medium sized models performed best on multiple datasets but showed no significant differences in WER from their smaller variants. Given the higher computational cost and training time of Medium models, Small Whisper models may be a more resource-efficient option for future clinical utilization in single-word speech tasks, especially if scaling to larger datasets.
Models fine-tuned on healthy (ft-h), patient (ft-p) and combined (ft-hp) SONIVA-Naming datasets showed comparable verbatim transcription and target detection performance on patient speech, which suggests that fine-tuning on age-matched healthy task-specific data may provide sufficient generalizability for patient data transcription for the same task. This is particularly relevant in clinical applications since collection of healthy speech data is more accessible, and can facilitate a larger dataset for fine-tuning.
Further, comparable performance between ft-h), (ft-p) and (ft-hp) models suggests that the main performance advantage comes from fine-tuning Whisper on short-form speech produced in the single-word Naming task, irrespective of the data source. This highlights the necessity of adapting foundation models to specific task paradigms to improve performance.
Similarly to the SONIVA-Naming test set, the fine-tuned Whisper models demonstrated significantly improved verbatim transcription performance on the unseen TORGO test set, compared to the baseline Whisper model, underscoring the limitations of using off-the-shelf pretrained models for out-of-domain clinical datasets and reinforcing the importance of model adaptation to short-form speech. However, the best-performing model—Medium ft-syn—still achieved an high Word Error Rate (WER) of 73.48%, which remains well above the threshold for clinical usability. Additionally, models fine-tuned on SONIVA patient data performed poorly when evaluated on TORGO, with the Medium model reaching a WER of 74.62%. This performance gap can be attributed to clinical heterogeneity: although both datasets capture speech and language-related impairments, SONIVA-Naming data focuses on stroke-induced aphasia that captures both motoric impairments and difficulties in linguistic processing, while TORGO assesses dysarthria and is only representative of motoric speech impairments. These findings highlight the necessity of tailoring pretrained models to the specific linguistic and pathological characteristics of each clinical condition to ensure robust generalization and clinical applicability.
In terms of target detection accuracy on the SONIVA-Naming task data, fine-tuned models showed higher True Positive and lower False Negative scores compared to baseline Whisper, making them more clinically applicable. In the case of patient data, the fine-tuned models produced more False Positive mistakes, as a result of failing to transcribe phonological errors made by patients (e.g., predicting target word ‘comb’ instead of ‘comg’). As Whisper is designed to predict the next most probable token in the sequence, these discrepancies are expected.
In addition to verbatim transcription, Whisper was used to complete a stroke-specific downstream task, aimed at estimating the level of speech impairment. In case of healthy data, the models initialized with ft-h encoder weights had the best F1 Macro in accuracy score prediction, suggesting the importance of fine-tuning. However, for patient data, models initialized with ft-p encoder weights improved performance on the Semantic and Phonology metrics but not on Dysfluency and Self-correction. Specifically, while the fine-tuned model outperformed previous studies that used Whisper encoder features for classifying dysfluencies in stuttering (F1 Macro of 0.71) [24], its performance was slightly lower than the baseline for Dysfluency. This suggests that while fine-tuning captures linguistic features like semantics and phonology, it may not capture more complex patterns in patient speech, such as hesitations and self-corrections, which are less vital for transcription but crucial for Dysfluency and Self-correction predictions, leading to worse performance on these metrics for patient speech. It is, however, worth noting that the absolute difference between the best baseline models and their fine-tuned version was negligible (1.45±0.06%).
In addition, the performance in the accuracy prediction task was higher on trials with unimpaired scores compared to impaired ones, especially in the healthy dataset. This is expected since in the case of healthy data, the models were exclusively fine-tuned on healthy speech where the number of impaired trials is inevitably limited, leading to smaller sample sizes and overall performance.
Despite this difference between model performance on impaired and unimpaired trials, the clinical validity analysis at the patient level confirmed that the predicted accuracy scores were clinically meaningful. Indeed, a significant difference between patients identified as impaired or unimpaired based on model predictions was observed in case of speech fluency, English proficiency status, and previous stroke history [34, 35, 36], which is expected. The results of the divergent validity analysis showed no differences in features unrelated to speech impairment such as cholesterol levels, sex and smoking status.
VI Limitations and Future Work
We note several limitations and areas for future work. First, patients were not stratified by speech impairment severity when split into training, validation, and test sets, which could lead to uneven distributions, affecting the model’s generalizability across impairment levels. However, this should not affect the main conclusions about Whisper drawn from this analysis given similar findings in the healthy-derived speech data.
In terms of the accuracy prediction task, the model performance could be improved by addressing the class imbalance issue by completing multiple binary classification tasks (instead of a multi-class one) and by using approaches other than downsampling. For instance, data perturbation and in-domain data augmentation could be applied [37]. Data perturbation alters audio characteristics like pitch, frequency, or tempo to create more samples [38], and has been used in previous studies for transcribing disordered and aphasic speech [39, 40]. In-domain data augmentation, instead, involves adding data from similar tasks from other datasets such as AphasiaBank [12].
In terms of transcription accuracy, the performance of Whisper with beam search rather than greedy decoding can be tested. The former may provide a more comprehensive conclusion on the applicability of Whisper for stroke-specific data. Greedy decoding consists of selecting the next token at each step in the output sequence, as the one with the highest probability [8, 41]. On the other hand, beam search is a decoding method that considers multiple possible tokens at each step, maintaining the most likely sequences and selecting the final output based on cumulative probability [41].
In order to obtain a broader overview on the application of foundation models to pathological speech in stroke, Whisper should be compared with other transformer-based architectures, such as Seamless M4T v2, a multimodal model capable of speech transcription that has shown competitive results compared to Whisper on ASR tasks [42]. Finally, future work should explore whether token-level loss analyses, encoder probing, or modality-specific pretraining can help disentangle the extent to which current models capture generalisable acoustic markers of clinical speech, particularly in settings where contextual compensation from the decoder is limited, such as single-word utterances.
VII Conclusion
In conclusion, in this study we evaluated the applicability of Whisper, a general foundation model, to a specific clinical case study on a single-word speech task in patients with stroke. Our findings confirmed that additional fine-tuning on single-word data is required to achieve efficient transcription performance that can facilitate clinically useful downstream tasks. Further, comparisons of fine-tuned model performance on in-domain and out-of-domain neurological disorder datasets suggests further need for model adaptation to disorder-specific data. To further assess the generalisability of the model, With such fine-tuning we can advance the automated analysis of pathological speech allowing for more efficient diagnosis and monitoring of stroke.
VIII Acknowledgment
The authors would like to thank the participants who took part in the IC3 study and the team from the University of Toronto for the TORGO database. We would like to acknowledge S.Brook and S.Moorjani for their contribution to data transcription. VG was supported by the Medical Research Council [MR/W00710X/1]. ZC was supported by the UK Engineering and Physical Sciences Research Council (EP/Y035216/1) for King’s Doctoral Training in DRIVE-Health and H2 Cognitive Designs Ltd. MD is employed as a Research Assistant by King’s College London and is supported by a Wellcome Trust grant (227099/Z/23/Z). The speech data was collected through support from the UK Medical Research Council (MR/T001402/1). Infrastructure support was provided by the NIHR Imperial Biomedical Research Centre and the NIHR Imperial Clinical Research Facility. The views expressed are those of the author(s) and not necessarily those of the NHS, the NIHR or the Department of Health and Social Care.